OpenStudy (jlastino):

How many gallons of paint are required to paint the walls of a building having l= 50 ft w= 40 ft. h =80ft. 1 gallon of paint can cover 500 square feet of the wall

OpenStudy (anonymous):

area of 1 wall = 50*40*80 sq ft gallons required = 50*40*80 / 500 = 50*32 / 5 = 320 gallons

OpenStudy (anonymous):

320

OpenStudy (jlastino):

So you don't need to find the surface area of the building?

OpenStudy (anonymous):

yes you do

OpenStudy (anonymous):

he did find the surface area

OpenStudy (anonymous):

actually i found the volume!!!!! - i just noticed that - it would only be a surface area if one of the numbers was the number of walls!! - so embarassing! can u check the question again jlastino!!

OpenStudy (jlastino):

haha.I think that was the volume. Anyways, i tried using the equation = 2lh +2wh

OpenStudy (anonymous):

for the area of the wall you just need w * h and then no. of walls in the building for total surface area

OpenStudy (jlastino):

(4wh/500) .Lol at my equation because 2lh would be the bases of the building (not the walls, of course). Thanks jimmy!

OpenStudy (anonymous):

yw